A jar contains 50 nickels and dimes worth $3.40. How many of each kind of coin are in the jar?
Answer by ewatrrr(24785)

 
Hi,
Let x represent number of dimes
Question states*** CENTS makes sense
10x + 5(50-x) = 340 cents
10x + 250-5x = 340 cents
5x = 90
x = 18, number of dimes. There are 32 Nickels
CHECKING our answer***
180 + 160 = 340 cents or $3.40
Wish You the Best in your Studies.